OpenStudy (anonymous):

y=-3/2=6 in standard form

OpenStudy (saifoo.khan):

You did some mistake typing the equation.

OpenStudy (saifoo.khan):

there are two equal signs.

OpenStudy (anonymous):

my bad its supposed to say y=-3/2+6

OpenStudy (saifoo.khan):

you missed the "x" as well?

OpenStudy (anonymous):

my bad .. y=-3/2x+6

OpenStudy (saifoo.khan):

y = mx + b Shift the mx to another side.

OpenStudy (anonymous):

your confusing me .. your

OpenStudy (anonymous):

your talking to a blonde

OpenStudy (saifoo.khan):

Lol. im trying to make you understand.. The slope form is, y=mx+b The only thing you change in standard form is shift the "mx" part to another side.

OpenStudy (anonymous):

y=mx+b has to convert to Ax+By=C

OpenStudy (saifoo.khan):

Yes. it's like, y = mx + b intercept form y - mx = b standard form

OpenStudy (anonymous):

mx is the slope

OpenStudy (saifoo.khan):

only m is the slope.

OpenStudy (saifoo.khan):

not x.

OpenStudy (anonymous):

eityou cant have a fraction as the answer

OpenStudy (saifoo.khan):